# Parametric Mean-Field Empirical Bayes in High-dimensional Linear Regression

### v1

@@ 3.1 Upper bound
To check consistency with parts (b) and (c) in 3.3, note that $h_{\theta}(t)=\operatorname{\mathbb{E}}\ell^{\prime}(\theta;B_{t,d_{0},\theta})=\operatorname{\mathbb{E}}[B_{t,d_{0},\theta}]-\theta=t$ is exactly linear. Thus, $h_{\theta}^{\prime\prime}\equiv 0$ for all $\theta$, and $\kappa\equiv 0$. In other words, every $\theta$ is a singular point in the sense of Remark 3.4. The limit in part (c) becomes degenerate, and does not reflect the tight order as the actual limit in (18).
@@ 3.2 Extension to misspecified priors
    $\kappa^{\star}(\theta):=\frac{1}{2}\operatorname{\mathbb{E}}[(\beta^{\star})^{2}]\operatorname{\mathbb{E}}[h_{\theta}^{\prime\prime}(W^{\star})].$
@@ 3.3 Lower bound
By the multivariate Cramér–Rao bound [e.g. see Theorem 2.6.1 in 35], for any unbiased estimator $\hat{\theta}_{\mathnormal{p}}$, its variance is bounded below by the inverse Fisher information:
@@
As the scaled limiting Fisher information matches the limiting variance in (15), we can conclude that the vEB estimator $\hat{\theta}^{\mathsf{vEB}}$ attains the optimal limiting variance when $p\ll n^{2/3}$. The proof of 3.5 builds upon the crucial observation that the Fisher information $\mathcal{I}_{p}(\theta_{0})$ can be expressed as moments of Random Field Ising Models (RFIMs, see A.1). Utilizing recent developments for RFIMs in [34], we establish moment bounds for the generalized summand $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}g(B_{i})$ (see Lemma A.8) for some smooth function $g(\cdot)$, which may be of independent interest.
@@ 4 Consequences of EB in downstream posterior inference
Here note that the EB posterior $\operatorname{\mathbb{P}}_{\hat{\theta}_{\mathnormal{p}}}$ depends on the estimated prior $\hat{\theta}_{\mathnormal{p}}$. So, the data $y$ is double-dipped to construct $\operatorname{\mathbb{P}}_{\hat{\theta}_{\mathnormal{p}}}$; once to write the likelihood $\operatorname{\mathbb{P}}(y\mid X,\beta)$ and second to estimate the prior $\hat{\theta}_{\mathnormal{p}}$. In this section, the notation $\hat{\theta}_{\mathnormal{p}}$ will denote any generic estimator, and $\hat{\theta}^{\mathsf{vEB}}$ will continue to denote the vEB estimator (11). Also, $\beta$ will denote a sample from the posterior (or its variational approximation), as opposed to the true regression coefficients $\beta^{\star}$ (see 3.3).
Our main goal is to understand whether the EB posterior is a good approximation of the oracle posterior. In this section, we answer this question in terms of posterior inference for (a) each coordinate, and (b) an one-dimensional projection. Due to technical reasons (mainly to use dependent concentration inequalities in [34]), this section will work under 3.5 and the asymptotic scaling $p=o(n^{2/3})$.
@@ 4.1 Estimating marginal distributions
Recall from 4.1 that $\operatorname{\mathbb{P}}_{\theta}$ denotes the usual posterior given a prior $\mu_{\theta}$. Define $\mathbb{Q}_{\theta}$ as the Mean-Field variational approximation of $\operatorname{\mathbb{P}}_{\theta}($:
@@
Here, note that $u_{i},\tau_{i}^{2}$ are oracle quantities that are functions of $\theta_{0}$, whereas $\hat{u}_{i},\hat{\tau}_{i}^{2}$ depend on the estimated hyperparameter $\hat{\theta}_{\mathnormal{p}}$. Also, note that these quantities are easily computable [see e.g. 34, 33], as opposed to the exact posterior mean and variance. In fact, these quantities connect to the Mean-Field optimizers from 2.2 as $u_{i}=u_{i,\theta_{0}},\hat{u}_{i}=u_{i,\hat{\theta}_{\mathnormal{p}}}$ (we will use the new notations for brevity).
@@
Suppose $p\ll n^{2/3}$ and assume Assumptions 3.1, 3.5. For each $i\leq p$, the following conclusions hold.
@@ 4.2 Uncertainty quantification for linear statistics
Now we move beyond pointwise estimation and discuss the consequences of vEB on uncertainty quantification. We first consider the oracle case. Due to our low SNR scaling where the posterior crucially depend on the choice of prior, one cannot establish a normal limit for each coordinate of the posterior [48]. Recently, jointly with Sumit Mukherjee, the authors established normal limiting distributions for one-dimensional projections of the posterior [33]. We restate part (a) of Theorem 3.1 and Proposition 3.2 in [33] below.
@@
In other words, under the oracle posterior, a CLT with a normal limiting distribution holds for the statistic $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}\beta_{i}$. Interestingly, the above result shows that one may replace the intractable mean and variance of the oracle posterior $\operatorname{\mathbb{P}}_{\theta_{0}}$ with the tractable counterparts that arise from the Mean-Field approximation $\mathbb{Q}_{\theta_{0}}$. This allows one to perform posterior inference, such as constructing credible intervals with explicit expressions. Note that establishing limit distributions of the oracle posterior requires the asymptotic requirement $p\ll n^{2/3}$.
@@
  • Can one modify the centering and variance in 4.2 to data-driven quantities (for example the plug-in counterparts using $\hat{\theta}_{\mathnormal{p}}$)?
The following result answers both questions, where the first question is true (see part (b)) but the plug-in part of the second question is generally false (see part (c)). Recalling the notation $B_{t,d,\theta}$ and $W_{\theta_{0}}$ from 3.1, define the following Jacobian $J:\textnormal{int}(\Theta)\to\mathbb{R}^{k}$:
@@
We elaborate on each conclusion. Part (a) extends the limit of the variational oracle posterior variance in (23) to the variance of the vEB posterior. Part (b) considers the EB posterior and shows that the oracle CLT in 4.2 can be extended by replacing $\theta_{0}$ with $\hat{\theta}_{\mathnormal{p}}$. Part (c) instead considers oracle posterior, and shows that the limit (25) extends when $\gamma=0$, or equivalently when $q$ is a contrast vector. However, for a general $\gamma\neq 0$, the limit no longer holds as the vEB posterior mean is biased. In other words, the estimation error of $\hat{\theta}_{\mathnormal{p}}$ results in a bias term with nontrivial fluctuations:
@@
This is why we have to marginalize the data $y$ to get a standard normal limit in (26). The notation $J(\theta)$ in (24) also arises from this bias, as we use a first-order Taylor expansion to approximate
@@
    has asymptotic $1-\alpha$ coverage under the EB posterior $\beta\sim{\operatorname{\mathbb{P}}}_{\hat{\theta}^{\mathsf{vEB}}}$:
@@
    has asymptotic $1-\alpha$ average coverage under the oracle posterior $\beta\sim\operatorname{\mathbb{P}}_{\theta_{0}}$:
@@
As noted in the introduction, this is in sharp contrast with the “merging” properties of the EB posterior in high SNR settings where the EB posterior is a higher-order approximation of the oracle posterior [c.f. Prop 3.1, 45]. In other words, the cited result states that the EB credible region can serve as an oracle credible region, without the need for variance inflation as in part (b).
@@ 5 Simulation studies
We conduct simulation studies to assess (i) the accuracy for recovering the prior hyperparameter $\theta_{0}$ and (ii) Bayesian uncertainty quantification for the regression coefficients $\beta$.
@@ 5.4 Improving the naive vEB estimator by debiasing
The following Figure 2 considers the one-dimensional symmetric Gaussian mixture prior $\mu_{\theta}\equiv\frac{1}{2}N(\theta,1)+\frac{1}{2}N(-\theta,1)$, and illustrates that bias correction leads to better finite sample performance, which is more apparent for larger $p$. Our debiasing formula arises from correcting the second-order expansion of the likelihood. More specifically, noting that the slower rate of convergence of $\hat{\theta}^{\mathsf{vEB}}$ arises from the second term in the RHS of (17), we correct for this bias by replacing $\kappa(\theta_{0})$ with $\kappa(\hat{\theta}^{\mathsf{vEB}})$. This is in contrast to the more common first-order debiasing in the frequentist literature for estimating high-dimensional regression coefficients $\beta$ [55, 4, 5, 37]. We conjecture that it may be possible to extend the limit (15) to higher dimensions $p\gg n^{3/4}$ by additionally correcting for higher-order error terms in the likelihood expansion (17).
@@ 6 Discussion
We discuss limitations of the current paper and avenues for future research.
@@
    Singular or hierarchical priors.  One common choice for practical empirical Bayes procedures is to use finite (Gaussian) mixtures to model the prior $\mu_{\theta}$. While our results allow such mixtures, we require standard regularity assumptions (see 3.2), which rules out singular models such as overfitted mixtures. Indeed, it is well-known that overfitted mixtures have a slower $p^{1/4}$-rate of convergence for parameter estimation, even under the non-hierarchical setting with $p$ i.i.d. observations [23]. It would be interesting to explore whether our phase transition for the vEB estimator persists under such singular priors. On a related note, it would be interesting to view mixture priors as another layer of hierarchy and consider inference for each mixture component (for example, suppose $\beta_{i}\mid Z_{i}\sim P_{Z_{i}}$ and $Z_{i}\sim\textsf{Categorical}(\pi)$, where $P_{1},\ldots,P_{M}$ denotes each mixture distribution, $Z_{i}$ denotes the mixture component, and $\pi$ is a $M$-dimensional weight vector). Related hierarchical modeling approaches were recently studied under the classical normal means EB problem [26] as well as oracle structured variational inference in regression [50], and it would be interesting to explore additional hierarchical modeling under our setup.
@@ Acknowledgments
The authors thank Sumit Mukherjee and Bodhisattva Sen for many helpful discussions throughout this work, and Zhou Fan for sharing the adaptive Langevin implementation in [21].
@@ Appendix A Proof of main theorems
    $\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\operatorname{\mathbb{E}}h(B_{w_{i},d_{i},\theta_{0}})\mid X$ $\operatorname*{\xrightarrow{\mathnormal{P}}}\operatorname{\mathbb{E}}h(B_{\theta_{0}}),\quad\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\big[\operatorname{\mathbb{E}}h(B_{w_{i},d_{i},\theta_{0}})\big]^{2}\mid X$ $\operatorname*{\xrightarrow{\mathnormal{P}}}\operatorname{\mathbb{E}}\Big[\operatorname{\mathbb{E}}\big(h(B_{\theta_{0}})\mid W_{\theta_{0}}\big)\Big]^{2}.$
@@
Finally, we state two technical moment-related lemmas using the sub-Gaussian assumption of the tilted prior (see R6 in 3.4). The following lemma provides convenient moment and cgf bounds under the sub-Gaussian assumption.
@@
$\sup_{\theta\in\Theta}\log\int e^{tb-\frac{db^{2}}{2}}d\mu_{\theta}(b)\lesssim 1+|t|^{4},\quad\operatorname{\mathbb{E}}|B_{t,d,\theta}|^{r}\lesssim_{r}1+|t|^{3r},\quad\forall r>0,~\theta\in\Theta,d>0.$
@@
    by [33, Equation (A.19)], define a simpler objective $M_{p}(\theta):=\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}g(w_{i},d_{0},\theta)$ to which we apply Lemma A.1. First note that the derivative of the map $t\mapsto g(t,d_{0},\theta)$ belong to $\operatorname{\mathsf{{Poly}}}(r)$ for all $\theta\in\Theta$, by Lemma A.5. Therefore, the first two conditions in Lemma A.1, LLN and unique maximizer, directly follow from Lemmas A.2 and A.3 respectively. For the third condition, it suffices to show the following bound:
@@
    $\sup_{\theta\in\Theta}\log\int e^{wb-\frac{d_{0}b^{2}}{2}}d\mu_{\theta}(b)\lesssim 1+w^{4},$
@@
    $=\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla F_{i}(\theta_{0})+\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla^{2}F_{i}(\theta_{0})(\hat{\theta}^{\mathsf{vEB}}-\theta_{0})+\frac{1}{2p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla^{3}F_{i}(\tilde{\theta}_{n})\times_{2}(\hat{\theta}^{\mathsf{vEB}}-\theta_{0})\times_{3}(\hat{\theta}^{\mathsf{vEB}}-\theta_{0})$
@@
    In the above expression, for a vector $v\in\mathbb{R}^{k}$, we have abbreviated $v^{k}$ as the $k$-fold outer product, i.e. $v^{2}:=v\otimes v,v^{3}:=v\otimes v\otimes v$. Recall that $B_{w_{i},d_{i},\theta}$ is defined in 3.1, and note that all expectations are conditional on $w_{i}$s. We claim that the third derivative term satisfies
@@
    $=\operatorname{\mathbb{E}}[\nabla^{2}\tilde{\ell}(\theta_{0};W_{\theta_{0}}),$
@@
  3. (iii)
    Finally, the individual conclusions in the theorem follow by using the limiting distribution of $p^{-1}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla F_{i}(\theta_{0})$ from Lemma A.4 alongside Slutsky’s theorem.
@@
$\mathcal{I}_{p}(\theta_{0})=-\operatorname{\mathbb{E}}\left[\frac{\partial^{2}\log m_{\theta}}{\partial\theta^{2}}\mid_{\theta=\theta_{0}}\right]=-\operatorname{\mathbb{E}}\left[\operatorname{\mathbb{E}}^{\tilde{\operatorname{\mathbb{P}}}^{A_{p},w}_{\theta_{0}}}\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla^{2}\ell(\theta_{0};B_{i})\Big]\right]-\operatorname{\mathbb{E}}\left[\operatorname{\textnormal{Var}}^{\tilde{\operatorname{\mathbb{P}}}^{A_{p},w}_{\theta_{0}}}\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla\ell(\theta_{0};B_{i})\Big]\right].$

### v2

@@ 3.1 Upper bound
To check consistency with parts (b) and (c) in 3.3, note that
$G_{\theta}(t)=\operatorname{\mathbb{E}}\ell^{\prime}(\theta;B_{t,d_{0},\theta})=\operatorname{\mathbb{E}}[B_{t,d_{0},\theta}]-\theta=t$
is exactly linear. Thus, $G_{\theta}^{\prime\prime}\equiv 0$ for all $\theta$, and $\kappa\equiv 0$. In other words, every $\theta$ is a singular point in the sense of Remark 3.4. The limit in part (c) becomes degenerate, and does not reflect the tight order as the actual limit in (19).
@@ 3.2 Extension to misspecified priors
    $\kappa^{\star}(\theta):=\frac{1}{2}\operatorname{\mathbb{E}}[(\beta^{\star})^{2}]\operatorname{\mathbb{E}}[G_{\theta}^{\prime\prime}(W^{\star})].$
@@ 3.3 Lower bound
By the multivariate Cramér–Rao bound [e.g. see Theorem 2.6.1 in 37], for any unbiased estimator $\hat{\theta}_{\mathnormal{p}}$, its variance is bounded below by the inverse Fisher information:
@@
As the scaled limiting Fisher information matches the limiting variance in (16), we can conclude that the vEB estimator $\hat{\theta}^{\mathsf{vEB}}$ attains the optimal limiting variance when $p\ll n^{2/3}$. The proof of 3.5 builds upon the crucial observation that the Fisher information $\mathcal{I}_{p}(\theta_{0})$ can be expressed as moments of Random Field Ising Models (RFIMs, see A.1). Utilizing recent developments for RFIMs in [36], we establish moment bounds for the generalized summand $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}g(\beta_{i})$ (see Lemma A.8) for some smooth function $g(\cdot)$, which may be of independent interest.
@@ 4 Consequences of EB in downstream posterior inference
Here note that the EB posterior $\operatorname{\mathbb{P}}_{\hat{\theta}_{\mathnormal{p}}}$ depends on the estimated prior $\hat{\theta}_{\mathnormal{p}}$. So, the data $y$ is double-dipped to construct $\operatorname{\mathbb{P}}_{\hat{\theta}_{\mathnormal{p}}}$; once to write the likelihood $\operatorname{\mathbb{P}}(y\mid X,\beta)$ and second to estimate the prior $\hat{\theta}_{\mathnormal{p}}$. Such notions of EB posteriors have been of interest for a wide range of statistical problems, including sparse linear regression [27, 66] as well as model selection [50].
In this section, the notation $\hat{\theta}_{\mathnormal{p}}$ will denote any generic estimator, and $\hat{\theta}^{\mathsf{vEB}}$ will continue to denote the vEB estimator (12). Also, $\beta$ will denote a sample from the posterior (or its variational approximation), as opposed to the true regression coefficients $\beta^{\star}$ (see 3.3).
Our main goal is to understand whether the EB posterior is a good approximation of the oracle posterior. In this section, we answer this question in terms of posterior inference for (a) each coordinate, and (b) a one-dimensional projection. Due to technical reasons (mainly to use dependent concentration inequalities in [36]), this section will work under 3.5 and the asymptotic scaling $p=o(n^{2/3})$.
@@ 4.1 Estimating marginal distributions
Recall from 4.1 that $\operatorname{\mathbb{P}}_{\theta}$ denotes the usual posterior given a prior $\mu_{\theta}$. Define $\mathbb{Q}_{\theta}$ as the Mean-Field variational approximation of $\operatorname{\mathbb{P}}_{\theta}$, i.e.,
@@
Here, note that $u_{i},\tau_{i}^{2}$ are oracle quantities that are functions of $\theta_{0}$, whereas $\hat{u}_{i},\hat{\tau}_{i}^{2}$ depend on the estimated hyperparameter $\hat{\theta}_{\mathnormal{p}}$. Also, note that these quantities are easily computable [see e.g. 36, 35], as opposed to the exact posterior mean and variance. In fact, these quantities connect to the Mean-Field optimizers from 2.2 as $u_{i}=u_{i,\theta_{0}},\hat{u}_{i}=u_{i,\hat{\theta}_{\mathnormal{p}}}$ (we will use the new notations for brevity).
@@
Suppose $p\ll n^{2/3}$ and assume Assumptions 3.1, 3.5. Consider any estimator $\hat{\theta}_{\mathnormal{p}}$. For each $i\leq p$, the following conclusions hold.
@@
    In particular, the coordinate-wise mean and variance satisfy the following:
@@ 4.2 Uncertainty quantification for linear statistics
Now we move beyond pointwise estimation and discuss the consequences of vEB on uncertainty quantification. We first consider the oracle case. Due to our low SNR scaling where the posterior crucially depends on the choice of prior, one cannot establish a normal limit for each coordinate of the posterior [11, 51]. Recently, jointly with Sumit Mukherjee, the authors established normal limiting distributions for one-dimensional projections of the posterior [35]. We restate part (a) of Theorem 3.1 and Proposition 3.2 in [35] below.
@@
In other words, under the oracle posterior, a CLT with a normal limiting distribution holds for the statistic $\sum_{\mathnormal{i}=1}^{\mathnormal{p}}q_{i}\beta_{i}$. Interestingly, the above result shows that one may replace the intractable mean and variance of the oracle posterior $\operatorname{\mathbb{P}}_{\theta_{0}}$ with the tractable counterparts that arise from the Mean-Field approximation $\mathbb{Q}_{\theta_{0}}$. This allows one to perform posterior inference, such as constructing credible intervals with explicit expressions. Note that establishing limit distributions of the oracle posterior requires the asymptotic scaling $p\ll n^{2/3}$.
@@
  • Can one modify the centering and variance in 4.2 to data-driven quantities (for example the plug-in counterparts using $\hat{\theta}_{\mathnormal{p}}$) to get a limit under the oracle posterior?
The following result answers both questions: the first holds (see part (b)), whereas the plug-in version of the second generally fails (see part (c)). Recalling the notation $B_{t,d,\theta}$ and $W_{\theta_{0}}$ from 3.1, define the following Jacobian $J:\textnormal{int}(\Theta)\to\mathbb{R}^{k}$:
@@
We elaborate on each conclusion. Part (a) extends the limit of the variational oracle posterior variance in (24) to the variance of the vEB posterior. Part (b) considers the EB posterior and shows that the oracle CLT in 4.2 can be extended by replacing $\theta_{0}$ with $\hat{\theta}_{\mathnormal{p}}$. Part (c) instead considers the oracle posterior, and shows that the limit (26) extends when $\gamma=0$, or equivalently when $q$ is a contrast vector. However, for a general $\gamma\neq 0$, the limit no longer holds as the vEB posterior mean is biased. In other words, the estimation error of $\hat{\theta}_{\mathnormal{p}}$ results in a bias term with nontrivial fluctuations:
@@
This is why we have to marginalize over $y$ to get a standard normal limit in (27). The notation $J(\theta)$ in (25) also arises from this bias, as we use a first-order Taylor expansion to approximate
@@
    has asymptotic $1-\alpha$ coverage under the EB posterior ${\operatorname{\mathbb{P}}}_{\hat{\theta}^{\mathsf{vEB}}}$:
@@
    has asymptotic $1-\alpha$ average coverage under the oracle posterior $\operatorname{\mathbb{P}}_{\theta_{0}}$:
@@
As noted in the introduction, this is in sharp contrast with the “merging” properties of the EB posterior in high SNR settings where the EB posterior is a higher-order approximation of the oracle posterior [c.f. Prop 3.1, 47]. In other words, the cited result states that the EB credible region can serve as an oracle credible region, without the need for variance inflation as in part (b).
@@ 5 Simulation studies
We conduct simulation studies to assess (i) the accuracy for recovering the prior hyperparameter $\theta_{0}$ and (ii) Bayesian uncertainty quantification for the regression coefficients $\beta$. After presenting the main simulation results in Sections 5.1-5.3, we also propose a way to improve the vEB estimator by debiasing in Section 5.4.
@@ 5.4 Improving the naive vEB estimator by debiasing
The following Figure 2 considers the one-dimensional symmetric Gaussian mixture prior $\mu_{\theta}\equiv\frac{1}{2}N(\theta,1)+\frac{1}{2}N(-\theta,1)$, and illustrates that bias correction leads to better finite sample performance, which is more apparent for larger $p$. Our debiasing formula arises from correcting the second-order expansion of the likelihood. More specifically, noting that the slower rate of convergence of $\hat{\theta}^{\mathsf{vEB}}$ arises from the second term in the RHS of (18), we correct for this bias by replacing $\kappa(\theta_{0})$ with $\kappa(\hat{\theta}^{\mathsf{vEB}})$. This is in contrast to the more common first-order debiasing in the frequentist literature for estimating high-dimensional regression coefficients $\beta$ [58, 4, 5, 39]. We conjecture that it may be possible to extend the limit (16) to higher dimensions $p\gg n^{3/4}$ by additionally correcting for higher-order error terms in (18).
@@ 6 Discussion
We discuss avenues for future research.
@@
    Singular or hierarchical priors.  One common choice for practical empirical Bayes procedures is to use finite (Gaussian) mixtures to model the prior $\mu_{\theta}$. While our results allow such mixtures, we require standard regularity assumptions (see 3.2), which rules out singular models such as overfitted mixtures. Indeed, it is well-known that overfitted mixtures have a slower $p^{1/4}$-rate of convergence for parameter estimation, even under the non-hierarchical setting with $p$ i.i.d. observations [24]. It would be interesting to explore whether our phase transition for the vEB estimator persists under such singular priors. On a related note, it would be interesting to view mixture priors as another layer of hierarchy and consider inference for each mixture component (for example, suppose $\beta_{i}\mid\Delta_{i}\sim P_{\Delta_{i}}$ and $\Delta_{i}\sim\textsf{Categorical}(\pi)$, where $P_{1},\ldots,P_{M}$ denotes each mixture distribution, $\Delta_{i}$ denotes the mixture component, and $\pi$ is a $M$-dimensional weight vector). Related hierarchical modeling approaches were recently studied under the classical normal means EB problem [28] as well as oracle structured variational inference in regression [53].
@@ Acknowledgments
The authors thank Sumit Mukherjee and Bodhisattva Sen for many helpful discussions throughout this work, Shunan Sheng and Bohan Wu for comments on the draft, and Zhou Fan for sharing the adaptive Langevin implementation in [22].
@@ Appendix A Proof of main theorems
    $\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\operatorname{\mathbb{E}}h(B_{w_{i},d_{i},\theta_{0}})\mid X$ $\operatorname*{\xrightarrow{\mathnormal{P}}}\operatorname{\mathbb{E}}h(B_{\theta_{0}}),~~\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\big[\operatorname{\mathbb{E}}h(B_{w_{i},d_{i},\theta_{0}})\big]^{2}\mid X$ $\operatorname*{\xrightarrow{\mathnormal{P}}}\operatorname{\mathbb{E}}\Big[\operatorname{\mathbb{E}}\big(h(B_{\theta_{0}})\mid W_{\theta_{0}}\big)\Big]^{2}.$
@@
Finally, we state three technical moment-related lemmas using the sub-Gaussian assumption of the tilted prior (see R6 in 3.4). The following lemma provides convenient moment and cgf bounds under the sub-Gaussian assumption.
@@
$\sup_{\theta\in\Theta}\log\int e^{tb-\frac{db^{2}}{2}}d\mu_{\theta}(b)\lesssim 1+|t|^{4},\quad\sup_{\theta\in\Theta}\operatorname{\mathbb{E}}|B_{t,d,\theta}|^{r}\lesssim_{r}1+|t|^{3r},\quad\forall r>0,~t\in\mathbb{R},~d>0.$
@@
    by [35, Equation (A.19)], define a simpler objective $M_{p}(\theta):=\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}g(w_{i},d_{0},\theta)$ to which we apply Lemma A.1. First note that the derivative of the map $t\mapsto g(t,d_{0},\theta)$ belong to $\operatorname{\mathsf{{Poly}}}(r)$ for all $\theta\in\Theta$, by Lemma A.5. Therefore, the first two conditions in Lemma A.1, LLN and unique maximizer, directly follow from Lemmas A.2 and A.3 respectively. For the third condition, it suffices to show the following bound:
@@
    $\sup_{\theta\in\Theta}\log\int e^{tb-\frac{d_{0}b^{2}}{2}}d\mu_{\theta}(b)\lesssim 1+t^{4},\quad\forall t.$
@@
    $=\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla F_{i}(\theta_{0})+\frac{1}{p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla^{2}F_{i}(\theta_{0})(\hat{\theta}^{\mathsf{vEB}}-\theta_{0})$
    $\quad+\frac{1}{2p}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla^{3}F_{i}(\tilde{\theta}_{n})\times_{2}(\hat{\theta}^{\mathsf{vEB}}-\theta_{0})\times_{3}(\hat{\theta}^{\mathsf{vEB}}-\theta_{0})$
@@
    Here, note that we use R2 to ensure that the derivatives in the RHS are well-defined. In the above expression, for a vector $v\in\mathbb{R}^{k}$, we have abbreviated $v^{k}$ as the $k$-fold outer product, i.e. $v^{2}:=v\otimes v,v^{3}:=v\otimes v\otimes v$. Recall that $B_{w_{i},d_{i},\theta}$ is defined in 3.1, and note that all expectations are conditional on $w_{i}$s. We claim that the third derivative term satisfies
@@
    $=\operatorname{\mathbb{E}}[\nabla^{2}\tilde{\ell}(\theta_{0};W_{\theta_{0}})],$
@@
Finally, the individual conclusions in the theorem follow by using the limiting distribution of $p^{-1}\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla F_{i}(\theta_{0})$ from Lemma A.4 alongside Slutsky’s theorem. ∎
@@
$\mathcal{I}_{p}(\theta_{0})$ $=-\operatorname{\mathbb{E}}\left[\frac{\partial^{2}\log m_{\theta}}{\partial\theta^{2}}\mid_{\theta=\theta_{0}}\right]$
$=-\operatorname{\mathbb{E}}\left[\operatorname{\mathbb{E}}^{\tilde{\operatorname{\mathbb{P}}}^{A_{p},w}_{\theta_{0}}}\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla^{2}\ell(\theta_{0};B_{i})\Big]\right]-\operatorname{\mathbb{E}}\left[\operatorname{\textnormal{Var}}^{\tilde{\operatorname{\mathbb{P}}}^{A_{p},w}_{\theta_{0}}}\Big[\sum_{\mathnormal{i}=1}^{\mathnormal{p}}\nabla\ell(\theta_{0};B_{i})\Big]\right].$